OpenStudy (anonymous):

How many solutions can be found for the equation 5x + 3(x − 1) = 10x − 2x − 3? None One Two Infinitely many

OpenStudy (anonymous):

@KendrickLamar2014

OpenStudy (anonymous):

@usernamenotfound

OpenStudy (jjcoolja12):

Let's solve your equation step-by-step. 5x+3(x−1)=10x−2x−3 Step 1: Simplify both sides of the equation. 5x+3(x−1)=10x−2x−3 Simplify: (Hide steps) 5x+(3)(x)+(3)(−1)=10x+−2x+−3(Distribute) 5x+3x+−3=10x+−2x+−3 (5x+3x)+(−3)=(10x+−2x)+(−3)(Combine Like Terms) 8x+−3=8x+−3 8x−3=8x−3 Step 2: Subtract 8x from both sides. 8x−3−8x=8x−3−8x −3=−3 Step 3: Add 3 to both sides. −3+3=−3+3 0=0 Answer: All real numbers are solutions.

OpenStudy (anonymous):

oh so its d?

OpenStudy (anonymous):

great thanks I have a few more questons could you help?

OpenStudy (jjcoolja12):

yea